# Tuning Vision-Language Models with Candidate Labels by Prompt Alignment

### v2

@@ 6 Conclusion
This work, for the first time, investigated the scenario when fine-tuning vision-language models (VLMs) with candidate labels. Throughout a series of experiments, we empirically demonstrated that prompt learning is superior to other fine-tuning methods when fine-tuning VLMs with candidate labels. However, as the ambiguity of candidate labels increases, the performance of prompt learning is degraded. To alleviate this issue, we proposed a framework that disambiguates candidate labels by aligning the dynamically mixed class posterior of the handcrafted and soft prompt with the model’s output to guide the learning process with candidate labels. Comprehensive experimental results on multiple benchmark datasets demonstrate the effectiveness of our proposed framework.

### v3

@@ 5 Conclusion
This work, for the first time, investigated the scenario when tuning vision-language models (VLMs) with candidate labels. Throughout a series of experiments, we empirically demonstrated that prompt learning combined with PLL training objectives in a vanilla way can learn from candidate labels. However, as the ambiguity of candidate labels increases, its performance is degraded. To alleviate this issue, we proposed a framework that enhances the robustness by aligning the dynamically mixed class posterior of the handcrafted and learnable prompt with the model’s output to guide the learning process with candidate labels. Comprehensive experimental results and analysis on multiple benchmark datasets demonstrate the effectiveness of our proposed framework.